OpenStudy (anonymous):

Solve this system of equations: 5x+4y=-8 2x-y+-11

OpenStudy (anonymous):

can you clarify the 2nd equation please

OpenStudy (anonymous):

Sorry about that here it goes. 5x+4y=-8 2x-y=-11

OpenStudy (anonymous):

okay, then what i would do is use substitution. Solve the second equation for y, you should get: y=11+2x. Now plug this value for y into the first equation. This will give you the x value. Finally plug the x value into either one of the equations and solve or y.

OpenStudy (anonymous):

Hmm... I got: x = -4 y = 3

OpenStudy (anonymous):

plug it in to the equation to ensure its correct

OpenStudy (anonymous):

ur correct